### v1

# Quantum-annealing-inspired algorithms for multijet clustering

Abstract
Jet clustering or reconstruction, a procedure to identify sprays of collimated particles originating from the fragmentation and hadronization of quarks and gluons, is a key component at high energy colliders. It is a complicated combinatorial optimization problem and requires intensive computing resources. In this study, we formulate jet reconstruction as a quadratic unconstrained binary optimization (QUBO) problem and introduce novel quantum-annealing-inspired algorithms for clustering multiple jets in electron-positron collision events. One of these quantum-annealing-inspired algorithms, ballistic simulated bifurcation, overcomes problems previously observed in multijet clustering with quantum-annealing approaches. We find that both the distance defined in the QUBO matrix as well as the prediction power of the QUBO solvers have crucial impacts on the multijet clustering performance. This study opens up a new approach to globally reconstruct multijet beyond dijet in one-go, in contrast to the traditional iterative method.
keywords:
High-energy physics , Jet clustering , Quantum computation , Quantum annealing , Simulated bifurcation
[…]
[first]organization=Institute of High Energy Physics, Chinese Academy of Sciences, state=Beijing, postcode=100049, country=China
[…]
1 Introduction
Jet clustering is a fundamental component at high energy colliders to determine the kinematics of the underlying processes governed by quantum chromodynamics (QCD). Due to color confinement, quarks and gluons produced by the collisions or the decays of heavy particles initiate sprays of collimated particles originating from their fragmentation and hadronization. Jets serve as reliable proxies to determine the original parton kinematics.
Jet reconstruction is a complicated combinatorial optimization problem and requires intensive computing resources. It dates back to the proposal by Sternman and Weinberg [1], and various algorithms have been developed over decades since then, as reviewed in Refs. [2, 3, 4, 5, 6]. The majority of widely used jet clustering algorithms are implemented in FastJet [7] for both hadron and $e^{+}e^{-}$ colliders and have successfully been used in various experiments including those at the Large Hadron Collider (LHC) [8, 9, 10, 11, 12].
As we face the unprecedented increase in luminosity at the High-Luminosity LHC (HL-LHC) [13] and future colliders under consideration such as the Circular Electron Positron Collider (CEPC) [14, 15, 16], new approaches are actively being investigated to overcome this challenge. Applications of quantum computing and algorithms have recently attracted much attention, and have been applied to track reconstruction for example [17, 18, 19, 20, 21, 22, 23, 24, 25]. It also naturally led to the recent investigations on the jet reconstruction using the quantum annealing (QA) [26, 27, 28, 29] and quantum gate machines [30, 31, 32].
In this work, we demonstrate the potential of quantum-annealing-inspired algorithms (QAIAs) to pursue multijet finding. In previous works using quantum annealing [26, 27, 28], reconstructing dijet has been pursued either with the thrust- or quantum-angle-based approach, with the latter providing higher performance. However, there was a significant degredation in performance in multijet events [28]. By replacing the quantum-angle with the $ee$-$k_{t}$ distance in the algorithm and using QAIAs, which will be detailed in Section 2, we succeed in maintaining or even slightly improving performance also for the multijet reconstruction.
2 Methodology
Jet clustering can be regarded as a combinatorial optimization problem, which could be formulated as a quadratic unconstrained binary optimization (QUBO) or Ising problem [33]. The difference between the two lies in whether to use zero/one binaries for the former or the $\pm$1 spins for the latter. The problem is designed in a way that the ground state of the QUBO/Ising model provides the correct answer. It is an NP (Nondeterministic Polynomial time) complete problem and the solution candidates diverge exponentially with the problem size. Quantum annealing (QA) machines by D-Wave based on the concept described in Ref. [34] and the coherent Ising machine (CIM) [35] are developed to efficienctly solve such kind of problems, for example. However, QA is generally known to provide suboptimal results when handling largely-connected graphs due to the limited connectivity of qubits and hardware noise [36, 37, 38, 39]. Such a trend is consistently observed in the previous jet clustering studies [26, 27].
The quantum bifurcation machine (QbM) [40, 41], a modified approach to CIM, carries out quantum adiabatic evolution with Kerr-nonlinear parametric oscillations. Simulated bifurcation [42] emulates the bifurcation phenomena with a classical nonlinear Hamiltonian, with some improvements introduced to the classical bifurcation machine (CbM), which was first proposed in parallel with QbM [40, 41]. This classical approach no longer has the foundation of quantum adiabatic theorem but can still find quasi-optimal solutions. Nevertheless, aSB is still prone to errors originating from the continuous treatment of spins in the differential equations. Two variants of SB are introduced to suppress such analog errors: ballistic SB (bSB) and discrete SB (dSB) [43]. bSB introduces complete inelastics walls to constrain the continuous values to be compatible with the spins, whereas the latter introduces partial discretization to the oscillator positions. The SB algorithms are found to be suitable for parallel computing and expected to outperform the CIM. The exact formulation of the SB algorithms is explained in A. In this study, bSB and dSB are considered. aSB is not considered as its performance is largely outperformed by bSB and dSB. The classical version of CIM (SimCIM) [44] showed visibly degraded performance, thus is also not considered. The SB algorithms used in this study are implemented in Huawei MindSpore Quantum [45, 46]. A simulated annealing library D-Wave Neal [47] is adopted as a benchmark for comparison to the SB algorithms.
2.1 Jet reconstruction as an Ising problem
As proposed in Refs. [26, 27, 28], jet reconstruction can be formulated in terms of a QUBO:
$O_{\textrm{QUBO}}({s_{i}})=\sum_{i,j=1}^{N_{\textrm{input}}}Q_{ij}s_{i}s_{j},$ (1)
where $s_{i}$ is the binary for each jet constituent to define which jet it is assigned to, $Q_{ij}$ is the QUBO matrix, which stores the distance between the $i$-th and $j$-th constituents, and $N_{\textrm{input}}$ is the number of inputs, namely the jet constituents. The thrust and quantum angle-based algorithms were considered in the previous studies. In this study, the quantum angle-based algorithm is considered as a benchmark and its QUBO is defined as:
$Q_{ij}=-\frac{\vec{p}_{i}\cdot\vec{p}_{j}}{2(|\vec{p}_{i}|\cdot|\vec{p}_{j}|)},$ (2)
where $\vec{p}_{i}$ is the momentum of the $i$-th jet constituent. It is compared to the QUBO based on the $ee$-$k_{t}$ distance [48]:
$Q_{ij}=2\textrm{min}(E_{i}^{2},E_{j}^{2})(1-\cos\theta_{ij}).$ (3)
where $E_{i}$ is the energy of the $i$-th jet constituent and $\theta_{ij}$ is the angle between the $i$-th and $j$-th jet constituents. The $ee$-$k_{t}$ or the so-called Durham algorithm [48] is the standard jet finding algorithm adopted in recent electron-positron colliders and is described in Section 2.2.
It is worth noting that the above formalism can only handle dijet clustering as is obvious from the binary implementation. In order to expand the method to multijet problems, the QUBO can be generalized to [26, 28]:
$O_{\textrm{QUBO}}^{\textrm{multijet}}({s_{i}^{(n)}})=\sum_{n=1}^{n_{\textrm{jet}}}\sum_{i,j=1}^{N_{\textrm{input}}}Q_{ij}s_{i}^{(n)}s_{j}^{(n)}+\lambda\sum_{i=1}^{N_{\textrm{input}}}\left(1-\sum_{n=1}^{n_{\textrm{jet}}}s_{i}^{(n)}\right)^{2},$ (4)
where $n$ considers the jet multiplicity and the binary $s_{i}^{(n)}$ is defined for each jet. The second term is introduced as the constraint to ensure that each jet constituent is assigned to a jet and only once. The coefficient of this penalty $\lambda$ must be taken large enough: $\lambda>N_{\textrm{input}}\textrm{max}_{ij}Q_{ij}$ [26]. In this study, the jet multiplicity is fixed to a certain value as is the case for the exclusive jet finding pursued at the electron positron colliders.
SB emulates quantum bifurcation phenomena within classical nonlinear Hamiltonian systems, thus the QUBO is converted to an Ising Hamiltonian as described in A.
2.2 Benchmark algorithm
The Durham, or $ee-k_{t}$ algorithm is adopted as our benchmark, since it is the standard jet finder at the electron positron colliders. It is implemented in the FastJet software package [7]. The algorithm computes the distance $d_{ij}$ between every pair of inputs $i$ and $j$:
$d_{ij}=2\textrm{min}(E_{i}^{2},E_{j}^{2})(1-\cos\theta_{ij}).$ (5)
It iteratively loops over and find the smallest $d_{ij}$ and recombine the two inputs into a single “particle”. For the exclusive mode, which we consider in this study, it terminates when the iteration reaches a fixed number of jets manually defined by the user. For the inclusive mode, usually adopted at hadron colliders, a scale $d_{\textrm{cut}}$ is defined by the user and the iteration stops when the minimum $d_{ij}$ exceeds the threshold. The latter is not considered in this study, since we only analyze the electron-positron collision events.
3 Dataset and event selection
We generate Monte Carlo (MC) simulated datasets for three physics processes, $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$, $e^{+}e^{-}\rightarrow ZH\rightarrow q\bar{q}b\bar{b}$ and $e^{+}e^{-}\rightarrow t\bar{t}\rightarrow bq\bar{q}^{\prime}\bar{b}\bar{q}^{\prime\prime}q^{\prime\prime\prime}$ at the center-of-mass energy of 91 GeV, 240 GeV and 350 GeV respectively, with MadGraph_aMC@NLO [49] for the matrix element calculation, Pythia8 [50] (v8.2, GPL-2) for the parton showering and hadronization, and Delphes [51] (v3.4.2, GPL-3) using the fourth detector concept [52] from the Circular Electron Positron Collider (CEPC), a future Higgs boson factory under consideration in China, for the fast simulation of detector effects. The Particle Flow (EFlow) objects are considered as the input for jet reconstruction. The three scenarios of the center-of-mass energy are in accordance with the CEPC proposal and are adopted here to evaluate the reconstruction performance for various jet multiplicities.
We only select events in which all the jets are within the detector acceptance, namely $\left|\cos~{}\theta\right|<0.9$. Furthermore, the separation of the two lowest leading jets:
$\sqrt{d_{nm}^{(\textrm{jet})}}=\sqrt{2\textrm{min}(E_{n}^{\textrm{(jet)}2},E_{m}^{\textrm{(jet)}2})(1-\cos\theta_{nm})},$ (6)
is required to be larger than 20 GeV. This selection significantly suppresses background with QCD (gluon) radiation with a minimal impact on the signal [53] and is adopted as a baseline pre-selection in this study. In the $ZH$ and $t\bar{t}$ events, we require them to have exactly two $b$-tagged jets. To simplify the analysis mentioned in Section 4, only the $ZH$ events with the $Z$ bosons decaying to non-$b$ quarks are generated in the above MC simulation sample.
4 Results
Figure 1: Event displays from a $t\bar{t}$ event with jet reconstructed by (a) $ee$-$k_{t}$ algorithm implemented in FastJet or with the $ee$-$k_{t}$-distanced QUBO approach using (b) bSB, (c) dSB or (d) D-Wave Neal. Each circle represents a particle flow candidate with its size proportional to the energy. Each color corresponds to an individual jet.
First, the predicted minimum energy by the three QAIAs are presented in Figure 5 for the three physics processes. For all cases, bSB outperforms dSB and D-Wave Neal, i.e. in terms of the mean predicted energy and stability against multiple measurements. The trend becomes more and more apparent as the physics process becomes more complicated; an order-of-magnitude improvement in the minimum energy prediction for the $t\bar{t}$ process using the $ee$-$k_{t}$ distance. dSB, on the other hand, does not succeed in predicting minimum energy and performance is roughly comparable with D-Wave Neal. It is worth noting that the QUBOs for jet reconstruction are generally fully-connected matrices unlike the sparse QUBOs defined in track reconstruction [17, 23]. QA is known for degraded performance in such a fully-connected case [27] and it is remarkable that bSB can still provide quasi-optimal solutions.
Figure 2: Constituent-matching efficiencies to FastJet for jets reconstructed by bSB (a,d,g), dSB (b,e,h) and D-Wave Neal (c,f,i) using $ee$-$k_{t}$ or angle-based QUBOs.
In order to evaluate the performance of jet clustering algorithms, the most naive approach would be to compare with the true assignment of constituents to a jet corresponding to the original parton. However, as also stated in Ref. [31], it is not possible to define such an assignment in a reasonable manner. The fundamental issue is that the final state hadrons are occasionally hadronized from quarks originating from different initial partons. The situation will be even more complicated and fundamentally ambiguous when high-order calculations come into play. Thus, it is impossible to define the “true” particle-parton association in a meaningful way.
Figure 3: Invariant masses of $Z$ bosons (a), Higgs bosons (b) and top quarks (c) with jets reconstructed by FastJet and QAIAs. Only FastJet and bSB are shown in (b,c) as dSB and Neal provide largely degraded performance and even fail to reconstruct some jets in the events.
Figure 4: Evolution of jet efficiencies against the running time from the three QAIAs presented for (a) $Z$ boson, (b) $ZH$, and (c) $t\bar{t}$ events. The $ee$-$k_{t}$ distance is adopted in the QUBO formulation. The solid lines are the average of multiple shots (100 for $Z$ and 50 for $ZH$ and $t\bar{t}$) and the envelopes represent the standard deviation from the mean of those shots.
We adopt a commonly taken approach [28, 31] to define a jet-constituent matching “efficiency” per jet by comparing to a corresponding classical algorithm, $ee$-$k_{t}$ implemented in FastJet, in this study, and quantify the percentage of the jet constituents clustered in the same way:
[…]
In the actual computation, the QAIAs return outputs of $n_{\textrm{jet}}$ arrays of jet constituent indices, where each array represents a jet. The jets reconstructed by FastJet are recorded in the same manner to compute the efficiency. The number of overlapping array components between FastJet and the QAIAs are used for the numerator in Eq. (7).
Figure 1 shows event displays from a $t\bar{t}$ event using the benchmark $ee$-$k_{t}$ algorithm implemented in FastJet as well as three QAIAs using the $ee$-$k_{t}$-distanced QUBO. Event displays with the angle-based QUBO as well as for the other two physics processes are presented in C. It is clearly seen that only bSB with the $ee$-$k_{t}$-distanced QUBO can successfully reconstruct jets for the three physics processes. Figure 2 shows the jet efficiency (Eq. 7), or resemblance to FastJet, to be more precise, evaluated for the three physics processes using the three QAIAs with two types of QUBOs: $ee$-$k_{t}$-based and angle-based. Several important findings can be seen from the figures. First, the $ee$-$k_{t}$-based QUBO outperforms the angle-based QUBO for all the physics processes. This is partially expected due to the consistent definition of the distance in the $ee$-$k_{t}$-based QUBO. More importantly, the angle-based QUBO stops to function for jet multiplicities beyond two; the jet efficiency largely degrades in the $ZH$ and $t\bar{t}$ events and even some jets have zero efficiency, meaning that they fail to be reconstructed despite the requirement of exclusive jet reconstruction. Most important of all, only bSB remains to succeed in jet reconstruction for all the three physics processes. With dSB and D-Wave Neal, the constituents are often chaotically and unreasonably assigned to a jet as can be seen in Figures 178 and 9. As stated above, fully-connected QUBOs are notoriously challenging to predict the minimum energy. bSB significantly outperforms dSB and Neal in this regard and demonstrates itself as a promising QUBO problem solver to handle multijet reconstruction.
As the benchmark $ee$-$k_{t}$ algorithm implemented in FastJet does not necessarily provide the exactly “correct” answer, the jet efficiency alone does not give us a fully decisive picture on the overall jet reconstruction performance. To evaluate the impact on the actual physics analysis, the invariant masses of the $Z$, Higgs bosons and top quarks are presented in Figure 3. In the $ZH$ events, the two $b$-tagged jets are assumed to originate from the Higgs bosons and are used in the mass reconstruction. The top-quark mass reconstruction is pursued in two steps as was done in Ref. [54]: the two $b$-jets are assumed to originate from the $b$-quarks and the light-flavor jet pairs with the least deviation from the $W$-boson mass $m_{W}$ are selected from the three possible permutations:
$\left|m_{ij}-m_{W}\right|+\left|m_{kl}-m_{W}\right|,$ (8)
where $i$, $j$, $k$, $l$ are the jet indices. Then one of the two possible combinations of these light-flavor jet pairs and the $b$-jets more compatible with the top-quark mass is adopted. We did not consider the $\chi^{2}$ method [55], kinematic likelihood [56] or state-of-the-art machine learning methods [57, 58, 59, 60, 61, 62] usually applied in hadron colliders, as the event topology and our assumption are simpler; the jet multiplicity is fixed to six and the $b$-jets are assumed to come from the $b$-quarks. In the $Z$ boson events, FastJet and bSB provide comparable performance, but dSB and D-Wave Neal already show visibly degraded mass resolution. In the $ZH$ and $t\bar{t}$ events, dSB and D-Wave Neal fail to reconstruct the jets in a reasonable manner as described above and are not shown in the figure. For these high-jet-multiplicity events, bSB starts to provide slightly better mass resolution than the baseline FastJet $ee$-$k_{t}$. This indicates that the global QUBO jet reconstruction using bSB may provide more precise clustering.
The execution time for each QAIA algorithm is evaluated on an AMD Ryzen 7 6800HS Creator Edition CPU and an NVIDIA A100 GPU. Figure 4 (6) presents the evolution of jet efficiencies (Ising energies) evaluated for the three QAIAs. We control the total runtime of the algorithms by setting different numbers of iterations. Only one CPU or GPU is used for fair comparison with D-Wave Neal. The average of the 100 (50) shots for the $Z$ boson ($ZH$ and $t\bar{t}$) events as well as the envelopes defined from the standard deviation from the mean are shown in the figure. bSB performs the best and rapidly converges to optimal values, whereas dSB and D-Wave Neal fails to reach reasonable jet efficiencies and Ising energies regardless of the running time.
5 Conclusion and Outlook
In our study, jet clustering is formulated as a QUBO problem and three QAIAs are adopted to pursue global reconstruction, which is a new approach compared to the traditional iterative reconstruction, implemented in FastJet, for example. The distance defined in the QUBO design has a significant impact on the reconstruction performance, particularly when the jet multiplicity is beyond two. It is found that the angle-based approach only provides reasonable performance for dijet events and alternative distances such as $ee$-$k_{t}$ are mandatory for higher jet multiplicities. For such multijet cases, it becomes exceedingly difficult even to approximately search for the “quasi”-ground state, ending up in a local minimum, which has an order-of-magnitude higher energy than the optimal states (Figure 5 (c,e)). Because of such a challenge, previous studies saw fairly degraded performance in multijet reconstruction. However, this study shows that a powerful QUBO problem solver, namely bSB among the three QAIAs, can find optimal solutions and demonstrates its outstanding capability to solve combinatorial optimization problems, despite the high QUBO connectivities. It is promising to see that this global jet reconstruction with bSB can bring in improvement to the invariant mass reconstruction in multijet events.
Being a “quantum-inspired” algorithm, bSB runs on classical computers and is suitable for parallel processing and usage of cutting-edge computing resources such as GPUs and FPGAs. It is important to note that we can flexibly balance speed and reconstruction precision in QAIAs. Namely, if precise energy resolution is not required, for low-level triggers, for example, we can run the reconstruction much faster. Thus, with the applicability to run on FPGAs, bSB may particularly provide an important option to be considered for triggers, during the $Z$-pole data taking at the CEPC, an “exa-byte”-level data taking comparable to the intensive HL-LHC conditions. The running time tends to be longer for high jet multiplicity events, and further investigations on the speed-up are ongoing.
Lastly, this study concentrates on $e^{+}e^{-}$ collider conditions and exclusive jet reconstruction, but an extension to inclusive jet reconstruction at hadron colliders would also be of interest and is left for future studies.
Acknowledgements
HO would like to thank Gang Li, Shudong Wang and Xu Gao for feedback regarding the CEPC 4th detector concept in Delphes. HO is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12075060 and the NSFC Basic Science Center Program for Joint Research on High Energy Frontier Particle Physics under Grant No. 12188102.
Appendix A Simulated bifurcation
SB emulates quantum bifurcation phenomena within classical nonlinear Hamiltonian systems. The Hamiltonian of an Ising model with $N$ spin $\{x_{i}\}_{i=1}^{N}\in\{-1,1\}^{N}$ is defined as
$H(x_{i})=\frac{1}{2}\sum_{ij}^{N}J_{ij}x_{i}x_{j}+\sum_{i}^{N}h_{i}x_{i},$ (9)
which is equivalent to Eq. (1) and they are related by:
$x_{i}$ $\longleftrightarrow 2s_{i}-1$ (10)
$J_{ij}$ $\longleftrightarrow\frac{Q_{ij}}{2}$ (11)
$h_{i}$ $\longleftrightarrow\frac{\sum_{j}Q_{ij}}{2}$ (12)
The first version of SB algorithms proposed, adiabatic SB (aSB) [42], is formulated as:
$\centering\dot{x}_{i}\@add@centering$ $=$ $\frac{\partial H_{\textrm{aSB}}}{\partial y_{i}}=a_{0}y_{i}$ (13)
$\dot{y}_{i}$ $=$ $-\frac{\partial H_{\textrm{aSB}}}{\partial x_{i}},$ (14)
$=$ $-\left[x_{i}^{2}+a_{0}-a(t)\right]x_{i}+c_{0}h_{i}+c_{0}\sum_{j=1}^{N}J_{ij}x_{j},$
$H_{\textrm{aSB}}$ $=$ $\frac{a_{0}}{2}\sum_{i=1}^{N}y_{i}^{2}+V_{\textrm{aSB}}$ (15)
$V_{\textrm{aSB}}$ $=$ $\sum_{i=1}^{N}\left(\frac{x_{i}^{4}}{4}+\frac{a_{0}-a(t)}{2}x_{i}^{2}\right)$ (16)
$-c_{0}\sum_{i=1}^{N}h_{i}x_{i}-\frac{c_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}x_{j},$
where $x_{i}$ and $y_{i}$ are respectively the position and momentum of the $i$-th Kerr-nonlinear oscillator corresponding to the $i$-th spin, $a_{0}$ and $c_{0}$ are positive constants (the detuning frequency for the former and the coupling strength for the latter), $a(t)$ is a time-dependent pumping amplitude that monotonically increases from zero to $a_{0}$, $H_{\textrm{aSB}}$ is the Hamiltonian, and $V_{\textrm{aSB}}$ is the potential energy in aSB.
The ballistic SB (bSB) algorithm [43] introduces complete inelastic scattering walls at $|x_{i}|>1$; i.e. $x_{i}$ is forced to be $x_{i}=\pm 1$ and $y_{i}=0$ when the particle passes the threshold:
$\centering\dot{x}_{i}\@add@centering$ $=$ $\frac{\partial H_{\textrm{bSB}}}{\partial y_{i}}=a_{0}y_{i}$ (17)
$\dot{y}_{i}$ $=$ $-\frac{\partial H_{\textrm{bSB}}}{\partial x_{i}},$ (18)
$=$ $-\left[a_{0}-a(t)\right]x_{i}+c_{0}h_{i}+c_{0}\sum_{j=1}^{N}J_{ij}x_{j},$
$H_{\textrm{bSB}}$ $=$ $\frac{a_{0}}{2}\sum_{i=1}^{N}y_{i}^{2}+V_{\textrm{aSB}}$ (19)
$V_{\textrm{bSB}}$ $=$ $\frac{a_{0}-a(t)}{2}\sum_{i=1}^{N}x_{i}^{2}-c_{0}\sum_{i=1}^{N}h_{i}x_{i}-\frac{c_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}x_{j}\quad$ (20)
$(\textrm{when}~{}|x_{i}|\leq 1~{}\textrm{for all}~{}x_{i}),$
$=$ $\infty\quad(\textrm{otherwise}).$
The quartic term is removed to avoid divergence, as the inelastic wall already plays similar roles to the nonlinear potential walls.
An alternative approach, discrete SB (dSB) [43], introduces discretization of the position, replacing $x_{i}$ with $\textrm{sgn}(x_{i})$:
$\centering\dot{x}_{i}\@add@centering$ $=$ $\frac{\partial H_{\textrm{dSB}}}{\partial y_{i}}=a_{0}y_{i}$ (21)
$\dot{y}_{i}$ $=$ $-\frac{\partial H_{\textrm{dSB}}}{\partial x_{i}},$ (22)
$=$ $-\left[a_{0}-a(t)\right]x_{i}+c_{0}h_{i}$
$+c_{0}\sum_{j=1}^{N}J_{ij}~{}\textrm{sgn}(x_{j}),$
$H_{\textrm{dSB}}$ $=$ $\frac{a_{0}}{2}\sum_{i=1}^{N}y_{i}^{2}+V_{\textrm{aSB}}$ (23)
$V_{\textrm{dSB}}$ $=$ $\frac{a_{0}-a(t)}{2}\sum_{i=1}^{N}x_{i}^{2}-c_{0}\sum_{i=1}^{N}h_{i}x_{i}$ (24)
$-\frac{c_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}~{}\textrm{sgn}(x_{j})\quad$
$(\textrm{when}~{}|x_{i}|\leq 1~{}\textrm{for all}~{}x_{i})$
$=$ $\infty\quad(\textrm{otherwise}),$
where the abovementioned partial discretization is introduced by replacing some $x_{i}$ with $\textrm{sgn}(x_{i})$.
Appendix B Ising energy predictions and evolution against running time
Figure 5 demonstrates the minimum Ising energies predicted by the three QAIAs for the three physics processes. bSB can achieve significantly improved predication of the minimum energies especially for high jet multiplicity events and QUBOs using the $ee$-$k_{t}$ distance. Figure 6 shows the evolution of Ising energies from the three QAIAs in the corresponding events.
Figure 5: Minimum Ising energies by the three QAIAs for a $Z$ boson (a,b), $ZH$ (c,d) and $t\bar{t}$ events (e,f) with the $ee$-$k_{t}$-based (a,c,e) or angle-based (b,d,f) distance considered in the QUBO formulation.
Figure 6: Evolution of Ising energies against the running time from the three QAIAs presented for (a) $Z$ boson, (b) $ZH$, and (c) $t\bar{t}$ events. The solid lines are the average of multiple shots (100 for $Z$ and 50 for $ZH$ and $t\bar{t}$) and the envelopes represent the standard deviation from the mean of those shots.
Appendix C Event displays
Event displays for the three physics processes are presented in Figures 7-9 reconstructed with the benchmark FastJet $ee$-$k_{t}$ algorithm or the bSB, dSB and D-Wave Neal using the $ee$-$k_{t}$ or the angled-based QUBOs.
Figure 7: Event displays from a Z event with jet reconstructed by (a) $ee$-$k_{t}$ algorithm implemented in FastJet or with the eekt- or angle-distanced QUBO approach respectively using (b,c) bSB, (d,e) dSB or (e,f) D-Wave Neal. Each circle represents a particle flow candidate with its size proportional to the energy. The same color corresponds to the same jet.
Figure 8: Event displays from a ZH event with jet reconstructed by (a) $ee$-$k_{t}$ algorithm implemented in FastJet or with the eekt- or angle-distanced QUBO approach respectively using (b,c) bSB, (d,e) dSB or (e,f) D-Wave Neal. Each circle represents a particle flow candidate with its size proportional to the energy. The same color corresponds to the same jet.
Figure 9: Event displays from a $t\bar{t}$ event with jet reconstructed by (a) $ee$-$k_{t}$ algorithm implemented in FastJet or with the eekt- or angle-distanced QUBO approach respectively using (b,c) bSB, (d,e) dSB or (e,f) D-Wave Neal. Each circle represents a particle flow candidate with its size proportional to the energy. The same color corresponds to the same jet.